OpenStudy (anonymous):

t-shirts for $12 each and sweatshirts for $15 each. You plan to spend $120 on t-shirts and sweatshirts. Write the equation that represents the situation.

OpenStudy (therealmeeeee):

12t + 15s = 120 Basically, substitute 0 for t, 0 for s, and then find the point where t = s (i.e. you spend as much on t-shirts as you do on sweatshirts.) 12(0) + 15s = 120 15s = 120 15s/15 = 120/15 s = 8 12t + 15(0) = 120 12t = 120 12t/12 =120/12 t = 10 12t + 15s = 120 12t = 15s 15s + 15s = 120 30s = 120 30s/30 = 120/30 s = 4 12t + 15(4) = 120 12t + 60 = 120 12t + 60 - 60 = 120 - 60 12t = 60 12t/12 = 60/12 t = 5 So you can by 8 sweatshirts, 10 t-shirts, or half and half (4 sweatshirts and 5 t-shirts).

OpenStudy (therealmeeeee):

did that help u

OpenStudy (anonymous):

Yeah thanks so much (:

OpenStudy (therealmeeeee):

click best reponse

OpenStudy (therealmeeeee):

ty

OpenStudy (anonymous):

np.. (;